Betsy strand regular hourly wage rate is $16 and she recieves an hourly rate of $24 for work in excess of 40 hours. During a January pay period, Betsy works 47 hours. Betsy Federal income tax witholding is $95 and she has no voluntary deductions. compute Betsy Strand gross earning earnings and net pay for the pay period. Assume that the Fica tax rate is 7.65%
Answer #1

Gross pay = Regular pay + Overtime pay

= (40*16) + (47-40)*24

= 808

Net pay = Gross pay - Federal income tax withheld - Fica taxes

= 808 - 95 - (808*7.65%)

= 808 - 95 - 61.81

= 651.19